OpenStudy (anonymous):

Using three consecutive odd integers, such that if 3 time the second is subtracted from the third, the difference equals 24. find the first integer

OpenStudy (mertsj):

x= first odd integer x+2 = second odd integer x+4=third odd integer x+4-3(x+2)=24 x+4-3x-6=24 -2x-2=24 -2x=26 x=-13 x+2=-11 x+4=-9